1. Suppose we have three random variables Y1 , Y2 , and Y3 .

Suppose we have three random variables Y, Y,, and Y,. The standard deviations of Y and Y, are both 3 and the standard deviation of Y is 2. The correlation coefficient between Y and Y, is-0.6. The covariance between Y and Y, is 0.5. Y is independent of Y 1. 1 2 a) (3 pts) Find Var(h + 3%) b) (3 pts) Find Cov(3h + 2⅓5½-%)
Answer #1

1.

a)

General formula for Variance of sum of two random variables X and Y is :

Var(aX+bY)=a^2Var(X)+b^2Var(Y)+2abCov(X,Y)

We need to find,

Var(Y_1+3Y_2) = Var(Y_1)+3^2Var(Y_2)+2*3*Cov(Y_1,Y_2)

We know that,

SD(Y_1) = 3 , SD(Y_2)=3, SD(Y_3)=2

So,

Var(A) = 32 = 9 : Var(Y) = 32-9. Var(⅓) = 22 = 4

We are also given,

Cor(Y_1,Y_2)=-0.6

Formula for correlation between two Random variables X and Y is :

Cov(X,Y) Cor(X,Y) = от Var(X) * Var(Y)

So,

Cor(Y_1,Y_2)=rac{Cov(Y_1,Y_2)}{sqrt{Var(Y_1)*Var(Y_2)}}=-0.6

Cor(y,%) 0.6 Var()Var.

Cor(y, Y)--0.6 * (VD * 9) =-5.4

So,

Var(Y_1+3Y_2) = Var(Y_1)+3^2Var(Y_2)+2*3*Cov(Y_1,Y_2)

9+32 * 9 2 * 3 * (-5.4)

57.6

b)

Now, we need to find,

Cov(3Y_1+2Y_3,5Y_2-Y_3)

= Cov(3Y_1,5Y_2)-Cov(3Y_1,Y_3)+Cov(2Y_3,5Y_2)-Cov(2Y_3,Y_3)

= 3*5*Cov(Y_1,Y_2)-3*Cov(Y_1,Y_3)+2*5*Cov(Y_3,Y_2)-2*Var(Y_3)

Since, Y2 and Y3 are independent , so ,

Cov(Y_2,Y_3)=0

and we also have,

Cov(Y_1,Y_2)=-5.4

Cov(Y_1,Y_3)=0.5

Putting all the values we get,

Cor(3Yİ + 2Y3.5y, _ Y3) 3 * 5 * (-5.4) _ 3 * (0.5) 2 * 5 * 0-2*4=-90.5